# Measuring Social Norms of Large Language Models

### v1

@@ Appendix D Summary of Skills
grade-12/choose-the-analysis-that-logically-connects-the-evidence-to-the-claim 54 76 100 100 84
grade-12/identify-run-on-sentences 50 56 45 81 82
grade-12/use-words-accurately-and-precisely 51 84 67 95 86
grade-12/avoid-double-illogical-and-unclear-comparisons 30 68 73 55 66
grade-12/choose-the-best-evidence-to-support-a-claim 46 23 100 72 100
grade-12/use-appeals-to-ethos-pathos-and-logos-in-persuasive-writing 50 47 27 26 78
grade-12/analyze-short-stories-set-2 40 68 74 61 82
@@
grade-12/choose-the-word-whose-connotation-and-denotation-best-match-the-sentence 42 100 100 100 100
grade-12/evaluate-counterclaims 38 42 100 55 100
grade-12/use-etymologies-to-determine-the-meanings-of-words 35 100 100 100 100
@@
grade-12/classify-the-figure-of-speech-euphemism-hyperbole-oxymoron-paradox 58 89 78 92 100
@@
grade-12/classify-the-figure-of-speech-anaphora-antithesis-apostrophe-assonance-chiasmus-understatement 73 100 100 100 100
grade-12/interpret-the-figure-of-speech 55 100 92 100 100
@@
Grade/Skill Random LLaMA2-70B-Chat SocialAgentLLaMA2-70B-Chat GPT3.5-Turbo SocialAgentGPT3.5-Turbo
grade-2/compare-urban-suburban-and-rural-areas 29 100 100 100 100
grade-2/rosa-parks 0 67 60 73 73
grade-3/thurgood-marshall 0 96 96 96 96
grade-3/rosa-parks 6 61 56 61 61
grade-3/the-bill-of-rights 56 100 100 100 100
grade-4/the-war-of-1812-events-and-effects 3 66 69 72 66
grade-4/understand-overall-supply-and-demand 33 100 100 100 100
grade-5/the-war-of-1812-causes 0 63 58 63 63
grade-5/the-war-of-1812-events-and-effects 0 66 66 72 72
grade-5/the-bill-of-rights 26 100 100 100 100
grade-5/understand-overall-supply-and-demand 14 100 100 100 100
grade-5/identify-shortage-and-surplus 29 40 30 50 100
grade-6/understand-overall-supply-and-demand 49 100 100 100 100
grade-7/the-bill-of-rights 31 100 100 100 100
grade-8/the-bill-of-rights 31 95 100 100 100
grade-8/understand-overall-supply-and-demand 40 100 100 100 100
grade-8/identify-primary-and-secondary-sources 37 75 75 75 75
grade-8/identify-facts-and-opinions 59 100 100 100 100
@@
Subject: Social Studies
Skill: the-american-revolution-new-british-taxes
Description: In the 1760s, smuggling was an important part of the colonial economy. What does smuggling mean?
@@
Choices: [(a) stealing goods from foreign merchants, (b) trading one good for another without the use of money, (c) shipping goods secretly or illegally, (d) producing metal from minerals and rocks]
Answer index: 2
Subject: Social Studies
Skill: world-war-i-the-road-to-war
Description: What was World War I originally called?
Choices: [(a) the First War, (b) the Great War, (c) the European Revolution, (d) European War I]
Answer index: 1
Subject: Social Studies
Skill: the-american-flag
@@
Description: What does each star on the American flag stand for?
Choices: [(a) a president, (b) a state, (c) a war, (d) a city]
Answer index: 1
Subject: Social Studies
Skill: identify-facts-and-opinions
Description: Which sentence states a fact?
@@
Choices: [(a) The Grand Canyon was carved by the Colorado River., (b) The Grand Canyon is the most breathtaking place the United States.]
Answer index: 0
Subject: Social Studies
Skill: state-government
Description: How are representatives in a state legislature chosen?
@@
Choices: [(a) Everyone in the state votes for all of the representatives., (b) Each district, or area, gets to choose one or two representatives., (c) The governor chooses all of the representatives., (d) The mayors of each city choose one or two representatives.]
Answer index: 1
Subject: Social Studies
Skill: the-empire-state-building
Description: Complete the sentence.
The Empire State Building was built during a tough time in American history. That time was known as ___.
Choices: [(a) the Colonial Period, (b) the Gold Rush, (c) the Reconstruction Era, (d) the Great Depression]
Answer index: 3
Subject: Social Studies
Skill: local-government
Description: Local governments serve which places?
Choices: [(a) countries, (b) towns, cities, and counties, (c) states]
Answer index: 1
Subject: Social Studies
Skill: the-american-revolution-the-rebellion-begins
Description: Complete the sentence.
At the First Continental Congress, the colonies agreed to ___ goods imported from Great Britain.
Choices: [(a) stop buying, (b) buy more, (c) destroy, (d) tax]
Answer index: 0
Subject: Social Studies
Skill: the-lincoln-memorial
Description: Complete the sentence.
Architect Henry Bacon designed the Lincoln Memorial in the early 1900s. He got his ideas from ___.
@@
Choices: [(a) architect Frank Lloyd Wright, (b) Gothic churches in France, (c) ancient Greek buildings, (d) President Warren G. Harding]
Answer index: 2
Subject: Social Studies
Skill: world-war-i-the-war-begins
Description: Complete the text.
There were two main sides in World War I. One side was called the ___ and the other was called the ___.
Choices: [(a) Axis powers . . . Allied powers, (b) Allied powers . . . Central powers, (c) Central powers . . . Axis powers, (d) Allied powers . . . Triple Entente]
Answer index: 1
Subject: Social Studies
Skill: world-war-i-the-road-to-peace
Description: Complete the text.
During the Paris Peace conference, one group of men was known as ___. They represented the most powerful countries at the conference.
Choices: [(a) the Three Emperors’ League, (b) the Axis Leaders, (c) the Big Four, (d) the enlightened despots]
Answer index: 2
Subject: Social Studies
Skill: the-statue-of-liberty
Description: The Statue of Liberty was finished in 1886. However, the builders’ goal had been to finish in 1876. Why was 1876 important?
@@
Choices: [(a) It was the year the Civil War ended., (b) It was the 100th anniversary of the Declaration of Independence., (c) New York City hosted the Olympics that year., (d) The president of France visited the United States for the first time that year.]
Answer index: 1
Subject: Social Studies
Skill: the-american-revolution-turning-the-tide-of-the-war
Description: The Continental Army suffered many losses in the autumn of 1776. Thomas Paine, a popular writer, was traveling with the army. In late December, he wrote an essay. Read the passage from Thomas Paine’s essay. Then answer the question below.
These are the times that try men’s souls. The summer soldier and the sunshine patriot will, in this crisis, shrink from the service of their country; but he that stands by it now, deserves the love and thanks of man and woman.
Which sentence best explains what he meant by these words?
Choices: [(a) He did not believe the Americans were doing the right thing., (b) General Washington did not know how to lead an army., (c) The British soldiers were afraid of the Americans., (d) The Continental Army was struggling, but it was fighting for a good cause.]
Answer index: 3
Subject: Social Studies
Skill: the-war-of-1812-causes
Description: The War of 1812 was between the United States and which other country?
Choices: [(a) France, (b) Canada, (c) Mexico, (d) Great Britain]
Answer index: 3
Subject: Social Studies
Skill: thanksgiving
Description: When is Thanksgiving celebrated in the United States?
@@
Choices: [(a) on December 5, (b) on January 12, (c) on the second Sunday of September, (d) on the fourth Thursday of November]
Answer index: 3
Subject: Social Studies
Skill: source-analysis-the-gettysburg-address
Description: The Gettysburg Address, one of the most famous documents in United States history, was written in 1863. At that time, Abraham Lincoln was president of the United States, and the Union and Confederate armies were fighting each other during the American Civil War.
Read the following facts about the Gettysburg Address. Then complete the sentence below.
Gettysburg is a town in Pennsylvania where an important Civil War battle was fought.
One of the meanings of the word address is a public speech.
Based on the information above, complete the sentence.
When people say, "President Abraham Lincoln gave the Gettysburg Address," they mean that President Lincoln ___.
Choices: [(a) lived at 1600 Pennsylvania Avenue, (b) wrote a book about an important Civil War battle, (c) made a speech after an important Civil War battle, (d) is buried in a graveyard in a town in Pennsylvania]
Answer index: 2
Subject: Social Studies
Skill: causes-of-the-american-revolution-the-french-and-indian-war
Description: Complete the text.
@@
Before the French and Indian War, there were about 20 times as many British colonists as French colonists. In the 1740s, both British and French colonists began to expand their claims on land in the Ohio River Valley. Both countries were interested in the valuable fur trade there, but ___ were more interested in settling the area.
Choices: [(a) The French, (b) the British]
Answer index: 1
Subject: Social Studies
Skill: the-civil-war-bull-run-to-the-emancipation-proclamation
@@
Description: In July 1861, a Union army marched south from Washington, D.C., into Virginia. Its goal was to seize Manassas Junction, an important railroad center. This plan led to the First Battle of Bull Run, the first major battle of the war.
baffled:confused
Complete the sentence.
Based on Whitman’s account, the First Battle of Bull Run was a ___ for the Union army.
Choices: [(a) defeat, (b) victory, (c) draw]
Answer index: 0
@@
Subject: Social Studies
Skill: identify-primary-and-secondary-sources
@@
Description: Imagine that you are a historian who wants to learn more about the fall of the Roman Empire in the 400s CE. You find the source described below. Read the description of the source. Then complete the sentence below.
The historian Edward Gibbon wrote the book The History of the Decline and Fall of the Roman Empire in 1776 CE.
Complete the sentence.
In this case, the book is a ___ source because it ___.
Choices: [(a) primary . . . is a firsthand account of the fall of the Roman Empire, (b) secondary . . . was written by a historian after the fall of the Roman Empire]
Answer index: 1
Subject: Social Studies
Skill: months-of-the-year
Description: What month comes right after February?
Choices: [(a) March, (b) December]
Answer index: 0
Subject: Social Studies
Skill: days-of-the-week
Description: What day comes right after Tuesday?
Choices: [(a) Saturday, (b) Friday, (c) Wednesday]
Answer index: 2
Subject: Social Studies
Skill: theodore-roosevelt
Description: Complete the sentence.
Theodore Roosevelt is best known as ___ from the early 1900s.
Choices: [(a) a president, (b) an artist, (c) a pilot, (d) an inventor]
Answer index: 0
Subject: Social Studies
Skill: george-washington-carver
@@
Description: George Washington Carver came up with how many ways to use peanuts?
Choices: [(a) around 3, (b) around 30, (c) around 300, (d) around 3,000]
Answer index: 2
Subject: Social Studies
Skill: cesar-chavez
Description: What made Cesar Chavez famous?
@@
Choices: [(a) He wrote many books about the Civil War., (b) He helped make life better for farm workers., (c) He invented a new way to bring water to crops., (d) He worked for women’s right to vote.]
Answer index: 1
Subject: Social Studies
Skill: john-deere
Description: Complete the sentence.
John Deere was a famous ___.
Choices: [(a) pilot, (b) soldier, (c) inventor, (d) writer]
Answer index: 2
Subject: Social Studies
Skill: pledge-of-allegiance
@@
Description: A pledge is a promise. Listen to this part of the Pledge of Allegiance. Pick the missing word.
I pledge ___ to the Flag of the United States of America . . .
Choices: [(a) allegiance, (b) honesty]
Answer index: 0
Subject: Social Studies
Skill: compare-urban-suburban-and-rural-areas
@@
Description: In which place would you be more likely to drive to see your neighbors?
Choices: [(a) an urban area, (b) a rural area]
Answer index: 1
Subject: Social Studies
Skill: abraham-lincoln
Description: Complete the sentence.
Abraham Lincoln grew up in a poor family. He got most of his education by ___.
@@
Choices: [(a) reading a lot of books, (b) taking college classes, (c) going to public school, (d) listening to the radio]
Answer index: 0
Subject: Social Studies
Skill: the-civil-war-the-first-battle-of-bull-run-to-gettysburg
Description: In 1862, the Union army in Washington, D.C., tried to capture the Confederacy’s capital of Richmond. Below is a letter from the Union army’s commander, General George B. McClellan, to President Lincoln’s Secretary of War.
@@
I have lost this battle because my force was too small. I have seen too many dead and wounded comrades to feel otherwise than this government has not sustained this army. If you do not do so now the game is lost. If I save this army now, I tell you plainly that I owe no thanks to you or any other person in Washington. You have done the best to sacrifice this army.
sustained: helped
Whom does McClellan blame for the Union army not capturing Richmond?
Choices: [(a) the Union soldiers, (b) himself, (c) the Confederate army, (d) the United States government]
Answer index: 3
Subject: Social Studies
Skill: new-england-colonies-religion-and-government
Description: Complete the text.
The Puritans were different from the Pilgrims who moved to New England in 1620. The Pilgrims wanted to ___ the Church of England, but the Puritans wanted to ___ it.
Choices: [(a) separate from . . . change, (b) change . . . separate from, (c) strengthen . . . destroy, (d) join . . . leave]
Answer index: 0
Subject: Social Studies
Skill: the-washington-monument
@@
Description: People started building the Washington Monument in 1848. It took 36 years to build. In what year was the Washington Monument completed?
Choices: [(a) 1854, (b) 1884, (c) 1898, (d) 1936]
Answer index: 1
Subject: Social Studies
Skill: the-civil-war-the-beginning-of-the-war
@@
Description: Following the Battle of Fort Sumter, President Lincoln called on loyal states to provide troops. How did Virginia, North Carolina, Tennessee, and Arkansas react to this call?
Choices: [(a) These states refused to send troops to either side., (b) These states joined the Confederacy., (c) These states demanded that the Confederate states surrender., (d) These states sent troops to the Union army.]
Answer index: 1
Subject: Social Studies
Skill: antebellum-period-slavery-and-politics-part-i
Description: Complete the text.
@@
By the 1820s, many Americans wanted to have an equal number of slave and free states. They wanted a balance of states so that the ___ would have the same number of members from each side.
Choices: [(a) Supreme Court, (b) Senate, (c) Executive Branch, (d) House of Representatives]
Answer index: 1
Subject: Social Studies
Skill: the-american-revolution-preparing-for-war
@@
Description: To oppose something means to be against it. Which one of these people strongly opposed American independence?
Choices: [(a) Thomas Paine, (b) John Adams, (c) King George III, (d) George Washington]
Answer index: 2
Subject: Social Studies
Skill: the-golden-gate-bridge
Description: Complete the sentence.
Leaders of the Golden Gate Bridge project made special rules to ___.
@@
Choices: [(a) protect birds in the area, (b) keep ferry boats in business, (c) use more recycled materials, (d) help keep workers safe]
Answer index: 3
Subject: Social Studies
Skill: identify-shortage-and-surplus
@@
Description: There are 300 concert tickets for sale. Tickets cost \$20 each. At that price, there are 200 people who want to buy a ticket.
Is there a surplus or a shortage of tickets?
Choices: [(a) surplus, (b) shortage]
Answer index: 0
Subject: Social Studies
Skill: the-new-deal
Description: Which important period in American history began in 1929?
Choices: [(a) the Civil War, (b) the Great Depression, (c) World War II, (d) the civil rights movement]
Answer index: 1
Subject: Social Studies
Skill: reconstruction
Description: In 1865, slavery was banned in the United States. How was it banned?
@@
Choices: [(a) a presidential veto, (b) a Supreme Court decision, (c) an amendment to the Constitution, (d) a national vote]
Answer index: 2
Subject: Social Studies
Skill: the-war-of-1812-events-and-effects
Description: Which country had a larger navy in 1812?
Choices: [(a) the United States, (b) Great Britain]
Answer index: 1
Subject: Social Studies
Skill: plymouth
@@
Description: In 1620, some of the Separatists made an agreement with a business in England called the Virginia Company. The Virginia Company would loan them money to buy supplies and hire a ship. Why did the Separatists need supplies and a ship?
Choices: [(a) to create a new country in Europe, (b) to find a trade route between Europe and Asia, (c) to build a new cathedral for the Church of England, (d) to start their own colony in North America]
Answer index: 3
@@
Subject: Social Studies
Skill: the-great-depression-part-i
Description: Who was the president of the United States at the start of the Great Depression?
Choices: [(a) William McKinley, (b) Herbert Hoover, (c) Harry Truman, (d) Gerald Ford]
Answer index: 1
Subject: Social Studies
Skill: the-american-revolution-struggle-for-independence
Description: At the Battle of Bunker Hill, one of the colonial officers gave the soldiers a famous order. Read the order. Then answer the question below.
Don’t fire until you can see the whites of their eyes!
What was one reason the officer gave this order?
Choices: [(a) He did not want the soldiers to waste any gunpowder or ammunition., (b) He wanted to make sure that the British soldiers were really going to attack., (c) He wanted to make sure that the soldiers were actually British., (d) He did not want the British soldiers to be surprised by the gunfire.]
Answer index: 0
Subject: Social Studies
Skill: federal-government
Description: Complete the sentence.
The three branches of government are the executive, legislative, and ___ branches.
Choices: [(a) federal, (b) defensive, (c) state, (d) judicial]
Answer index: 3
Subject: Social Studies
Skill: the-bill-of-rights
Description: Complete the sentences.

### v2

@@ Appendix D Summary of Skills
grade-12/choose-the-analysis-that-logically-connects-the-evidence-to-the-claim
54
76
100
100
84
grade-12/identify-run-on-sentences
50
56
45
81
82
grade-12/use-words-accurately-and-precisely
51
84
67
95
86
grade-12/avoid-double-illogical-and-unclear-comparisons
30
68
73
55
66
grade-12/choose-the-best-evidence-to-support-a-claim
46
23
100
72
100
grade-12/use-appeals-to-ethos-pathos-and-logos-in-persuasive-writing
50
47
27
26
78
grade-12/analyze-short-stories-set-2
40
68
74
61
82
@@
grade-12/choose-the-word-whose-connotation-and-denotation-best-match-the-sentence
42
100
100
100
100
grade-12/evaluate-counterclaims
38
42
100
55
100
grade-12/use-etymologies-to-determine-the-meanings-of-words
35
100
100
100
100
@@
grade-12/classify-the-figure-of-speech-euphemism-hyperbole-oxymoron-paradox
58
89
78
92
100
@@
grade-12/classify-the-figure-of-speech-anaphora-antithesis-apostrophe-assonance-chiasmus-understatement
73
100
100
100
100
grade-12/interpret-the-figure-of-speech
55
100
92
100
100
@@
Grade/Skill
Random
LLaMA2-70B-Chat
[…]
GPT3.5-Turbo
[…]
grade-2/compare-urban-suburban-and-rural-areas
29
100
100
100
100
grade-2/rosa-parks
0
67
60
73
73
grade-3/thurgood-marshall
0
96
96
96
96
grade-3/rosa-parks
6
61
56
61
61
grade-3/the-bill-of-rights
56
100
100
100
100
grade-4/the-war-of-1812-events-and-effects
3
66
69
72
66
grade-4/understand-overall-supply-and-demand
33
100
100
100
100
grade-5/the-war-of-1812-causes
0
63
58
63
63
grade-5/the-war-of-1812-events-and-effects
0
66
66
72
72
grade-5/the-bill-of-rights
26
100
100
100
100
grade-5/understand-overall-supply-and-demand
14
100
100
100
100
grade-5/identify-shortage-and-surplus
29
40
30
50
100
grade-6/understand-overall-supply-and-demand
49
100
100
100
100
grade-7/the-bill-of-rights
31
100
100
100
100
grade-8/the-bill-of-rights
31
95
100
100
100
grade-8/understand-overall-supply-and-demand
40
100
100
100
100
grade-8/identify-primary-and-secondary-sources
37
75
75
75
75
grade-8/identify-facts-and-opinions
59
100
100
100
100
@@
Subject: Social Studies Skill: the-american-revolution-new-british-taxes Description: In the 1760s, smuggling was an important part of the colonial economy. What does smuggling mean? Choices: [(a) stealing goods from foreign merchants, (b) trading one good for another without the use of money, (c) shipping goods secretly or illegally, (d) producing metal from minerals and rocks] Answer index: 2
Subject: Social Studies Skill: world-war-i-the-road-to-war Description: What was World War I originally called? Choices: [(a) the First War, (b) the Great War, (c) the European Revolution, (d) European War I] Answer index: 1
@@
Subject: Social Studies Skill: the-american-flag Description: What does each star on the American flag stand for? Choices: [(a) a president, (b) a state, (c) a war, (d) a city] Answer index: 1
@@
Subject: Social Studies Skill: identify-facts-and-opinions Description: Which sentence states a fact? Choices: [(a) The Grand Canyon was carved by the Colorado River., (b) The Grand Canyon is the most breathtaking place the United States.] Answer index: 0
@@
Subject: Social Studies Skill: state-government Description: How are representatives in a state legislature chosen? Choices: [(a) Everyone in the state votes for all of the representatives., (b) Each district, or area, gets to choose one or two representatives., (c) The governor chooses all of the representatives., (d) The mayors of each city choose one or two representatives.] Answer index: 1
Subject: Social Studies Skill: the-empire-state-building Description: Complete the sentence. The Empire State Building was built during a tough time in American history. That time was known as ___. Choices: [(a) the Colonial Period, (b) the Gold Rush, (c) the Reconstruction Era, (d) the Great Depression] Answer index: 3
Subject: Social Studies Skill: local-government Description: Local governments serve which places? Choices: [(a) countries, (b) towns, cities, and counties, (c) states] Answer index: 1
Subject: Social Studies Skill: the-american-revolution-the-rebellion-begins Description: Complete the sentence. At the First Continental Congress, the colonies agreed to ___ goods imported from Great Britain. Choices: [(a) stop buying, (b) buy more, (c) destroy, (d) tax] Answer index: 0
@@
Subject: Social Studies Skill: the-lincoln-memorial Description: Complete the sentence. Architect Henry Bacon designed the Lincoln Memorial in the early 1900s. He got his ideas from ___. Choices: [(a) architect Frank Lloyd Wright, (b) Gothic churches in France, (c) ancient Greek buildings, (d) President Warren G. Harding] Answer index: 2
Subject: Social Studies Skill: world-war-i-the-war-begins Description: Complete the text. There were two main sides in World War I. One side was called the ___ and the other was called the ___. Choices: [(a) Axis powers . . . Allied powers, (b) Allied powers . . . Central powers, (c) Central powers . . . Axis powers, (d) Allied powers . . . Triple Entente] Answer index: 1
Subject: Social Studies Skill: world-war-i-the-road-to-peace Description: Complete the text. During the Paris Peace conference, one group of men was known as ___. They represented the most powerful countries at the conference. Choices: [(a) the Three Emperors’ League, (b) the Axis Leaders, (c) the Big Four, (d) the enlightened despots] Answer index: 2
@@
Subject: Social Studies Skill: the-statue-of-liberty Description: The Statue of Liberty was finished in 1886. However, the builders’ goal had been to finish in 1876. Why was 1876 important? Choices: [(a) It was the year the Civil War ended., (b) It was the 100th anniversary of the Declaration of Independence., (c) New York City hosted the Olympics that year., (d) The president of France visited the United States for the first time that year.] Answer index: 1
Subject: Social Studies Skill: the-american-revolution-turning-the-tide-of-the-war Description: The Continental Army suffered many losses in the autumn of 1776. Thomas Paine, a popular writer, was traveling with the army. In late December, he wrote an essay. Read the passage from Thomas Paine’s essay. Then answer the question below. These are the times that try men’s souls. The summer soldier and the sunshine patriot will, in this crisis, shrink from the service of their country; but he that stands by it now, deserves the love and thanks of man and woman. Which sentence best explains what he meant by these words? Choices: [(a) He did not believe the Americans were doing the right thing., (b) General Washington did not know how to lead an army., (c) The British soldiers were afraid of the Americans., (d) The Continental Army was struggling, but it was fighting for a good cause.] Answer index: 3
Subject: Social Studies Skill: the-war-of-1812-causes Description: The War of 1812 was between the United States and which other country? Choices: [(a) France, (b) Canada, (c) Mexico, (d) Great Britain] Answer index: 3
@@
Subject: Social Studies Skill: thanksgiving Description: When is Thanksgiving celebrated in the United States? Choices: [(a) on December 5, (b) on January 12, (c) on the second Sunday of September, (d) on the fourth Thursday of November] Answer index: 3
Subject: Social Studies Skill: source-analysis-the-gettysburg-address Description: The Gettysburg Address, one of the most famous documents in United States history, was written in 1863. At that time, Abraham Lincoln was president of the United States, and the Union and Confederate armies were fighting each other during the American Civil War. Read the following facts about the Gettysburg Address. Then complete the sentence below. Gettysburg is a town in Pennsylvania where an important Civil War battle was fought. One of the meanings of the word address is a public speech. Based on the information above, complete the sentence. When people say, "President Abraham Lincoln gave the Gettysburg Address," they mean that President Lincoln ___. Choices: [(a) lived at 1600 Pennsylvania Avenue, (b) wrote a book about an important Civil War battle, (c) made a speech after an important Civil War battle, (d) is buried in a graveyard in a town in Pennsylvania] Answer index: 2
@@
Subject: Social Studies Skill: causes-of-the-american-revolution-the-french-and-indian-war Description: Complete the text. Before the French and Indian War, there were about 20 times as many British colonists as French colonists. In the 1740s, both British and French colonists began to expand their claims on land in the Ohio River Valley. Both countries were interested in the valuable fur trade there, but ___ were more interested in settling the area. Choices: [(a) The French, (b) the British] Answer index: 1
@@
Subject: Social Studies Skill: the-civil-war-bull-run-to-the-emancipation-proclamation Description: In July 1861, a Union army marched south from Washington, D.C., into Virginia. Its goal was to seize Manassas Junction, an important railroad center. This plan led to the First Battle of Bull Run, the first major battle of the war. baffled:confused Complete the sentence. Based on Whitman’s account, the First Battle of Bull Run was a ___ for the Union army. Choices: [(a) defeat, (b) victory, (c) draw] Answer index: 0
@@
Subject: Social Studies Skill: identify-primary-and-secondary-sources Description: Imagine that you are a historian who wants to learn more about the fall of the Roman Empire in the 400s CE. You find the source described below. Read the description of the source. Then complete the sentence below. The historian Edward Gibbon wrote the book The History of the Decline and Fall of the Roman Empire in 1776 CE. Complete the sentence. In this case, the book is a ___ source because it ___. Choices: [(a) primary . . . is a firsthand account of the fall of the Roman Empire, (b) secondary . . . was written by a historian after the fall of the Roman Empire] Answer index: 1
Subject: Social Studies Skill: months-of-the-year Description: What month comes right after February? Choices: [(a) March, (b) December] Answer index: 0
Subject: Social Studies Skill: days-of-the-week Description: What day comes right after Tuesday? Choices: [(a) Saturday, (b) Friday, (c) Wednesday] Answer index: 2
Subject: Social Studies Skill: theodore-roosevelt Description: Complete the sentence. Theodore Roosevelt is best known as ___ from the early 1900s. Choices: [(a) a president, (b) an artist, (c) a pilot, (d) an inventor] Answer index: 0
@@
Subject: Social Studies Skill: george-washington-carver Description: George Washington Carver came up with how many ways to use peanuts? Choices: [(a) around 3, (b) around 30, (c) around 300, (d) around 3,000] Answer index: 2
@@
Subject: Social Studies Skill: cesar-chavez Description: What made Cesar Chavez famous? Choices: [(a) He wrote many books about the Civil War., (b) He helped make life better for farm workers., (c) He invented a new way to bring water to crops., (d) He worked for women’s right to vote.] Answer index: 1
Subject: Social Studies Skill: john-deere Description: Complete the sentence. John Deere was a famous ___. Choices: [(a) pilot, (b) soldier, (c) inventor, (d) writer] Answer index: 2
@@
Subject: Social Studies Skill: pledge-of-allegiance Description: A pledge is a promise. Listen to this part of the Pledge of Allegiance. Pick the missing word. I pledge ___ to the Flag of the United States of America . . . Choices: [(a) allegiance, (b) honesty] Answer index: 0
@@
Subject: Social Studies Skill: compare-urban-suburban-and-rural-areas Description: In which place would you be more likely to drive to see your neighbors? Choices: [(a) an urban area, (b) a rural area] Answer index: 1
@@
Subject: Social Studies Skill: abraham-lincoln Description: Complete the sentence. Abraham Lincoln grew up in a poor family. He got most of his education by ___. Choices: [(a) reading a lot of books, (b) taking college classes, (c) going to public school, (d) listening to the radio] Answer index: 0
@@
Subject: Social Studies Skill: the-civil-war-the-first-battle-of-bull-run-to-gettysburg Description: In 1862, the Union army in Washington, D.C., tried to capture the Confederacy’s capital of Richmond. Below is a letter from the Union army’s commander, General George B. McClellan, to President Lincoln’s Secretary of War. I have lost this battle because my force was too small. I have seen too many dead and wounded comrades to feel otherwise than this government has not sustained this army. If you do not do so now the game is lost. If I save this army now, I tell you plainly that I owe no thanks to you or any other person in Washington. You have done the best to sacrifice this army. sustained: helped Whom does McClellan blame for the Union army not capturing Richmond? Choices: [(a) the Union soldiers, (b) himself, (c) the Confederate army, (d) the United States government] Answer index: 3
Subject: Social Studies Skill: new-england-colonies-religion-and-government Description: Complete the text. The Puritans were different from the Pilgrims who moved to New England in 1620. The Pilgrims wanted to ___ the Church of England, but the Puritans wanted to ___ it. Choices: [(a) separate from . . . change, (b) change . . . separate from, (c) strengthen . . . destroy, (d) join . . . leave] Answer index: 0
@@
Subject: Social Studies Skill: the-washington-monument Description: People started building the Washington Monument in 1848. It took 36 years to build. In what year was the Washington Monument completed? Choices: [(a) 1854, (b) 1884, (c) 1898, (d) 1936] Answer index: 1
@@
Subject: Social Studies Skill: the-civil-war-the-beginning-of-the-war Description: Following the Battle of Fort Sumter, President Lincoln called on loyal states to provide troops. How did Virginia, North Carolina, Tennessee, and Arkansas react to this call? Choices: [(a) These states refused to send troops to either side., (b) These states joined the Confederacy., (c) These states demanded that the Confederate states surrender., (d) These states sent troops to the Union army.] Answer index: 1
@@
Subject: Social Studies Skill: antebellum-period-slavery-and-politics-part-i Description: Complete the text. By the 1820s, many Americans wanted to have an equal number of slave and free states. They wanted a balance of states so that the ___ would have the same number of members from each side. Choices: [(a) Supreme Court, (b) Senate, (c) Executive Branch, (d) House of Representatives] Answer index: 1
@@
Subject: Social Studies Skill: the-american-revolution-preparing-for-war Description: To oppose something means to be against it. Which one of these people strongly opposed American independence? Choices: [(a) Thomas Paine, (b) John Adams, (c) King George III, (d) George Washington] Answer index: 2
@@
Subject: Social Studies Skill: the-golden-gate-bridge Description: Complete the sentence. Leaders of the Golden Gate Bridge project made special rules to ___. Choices: [(a) protect birds in the area, (b) keep ferry boats in business, (c) use more recycled materials, (d) help keep workers safe] Answer index: 3
@@
Subject: Social Studies Skill: identify-shortage-and-surplus Description: There are 300 concert tickets for sale. Tickets cost \$20 each. At that price, there are 200 people who want to buy a ticket. Is there a surplus or a shortage of tickets? Choices: [(a) surplus, (b) shortage] Answer index: 0
Subject: Social Studies Skill: the-new-deal Description: Which important period in American history began in 1929? Choices: [(a) the Civil War, (b) the Great Depression, (c) World War II, (d) the civil rights movement] Answer index: 1
@@
Subject: Social Studies Skill: reconstruction Description: In 1865, slavery was banned in the United States. How was it banned? Choices: [(a) a presidential veto, (b) a Supreme Court decision, (c) an amendment to the Constitution, (d) a national vote] Answer index: 2
Subject: Social Studies Skill: the-war-of-1812-events-and-effects Description: Which country had a larger navy in 1812? Choices: [(a) the United States, (b) Great Britain] Answer index: 1
@@
Subject: Social Studies Skill: plymouth Description: In 1620, some of the Separatists made an agreement with a business in England called the Virginia Company. The Virginia Company would loan them money to buy supplies and hire a ship. Why did the Separatists need supplies and a ship? Choices: [(a) to create a new country in Europe, (b) to find a trade route between Europe and Asia, (c) to build a new cathedral for the Church of England, (d) to start their own colony in North America] Answer index: 3
@@
Subject: Social Studies Skill: the-great-depression-part-i Description: Who was the president of the United States at the start of the Great Depression? Choices: [(a) William McKinley, (b) Herbert Hoover, (c) Harry Truman, (d) Gerald Ford] Answer index: 1
Subject: Social Studies Skill: the-american-revolution-struggle-for-independence Description: At the Battle of Bunker Hill, one of the colonial officers gave the soldiers a famous order. Read the order. Then answer the question below. Don’t fire until you can see the whites of their eyes! What was one reason the officer gave this order? Choices: [(a) He did not want the soldiers to waste any gunpowder or ammunition., (b) He wanted to make sure that the British soldiers were really going to attack., (c) He wanted to make sure that the soldiers were actually British., (d) He did not want the British soldiers to be surprised by the gunfire.] Answer index: 0
Subject: Social Studies Skill: federal-government Description: Complete the sentence. The three branches of government are the executive, legislative, and ___ branches. Choices: [(a) federal, (b) defensive, (c) state, (d) judicial] Answer index: 3